# Wetting on Silicone Surfaces

### v1

@@ I Introduction
Silicones are abundant materials with global production margins exceeding a million tonnes annually. As a surface coating, the material is widely used in numerous industries, ranging from the medical and energy sectors to the personal care and automotive/aerospace industries. The low surface energy of silicones promotes low friction and adhesion with contacting solids and liquids, providing desirable properties for self-cleaning, lubrication, anti-icing/-biofouling, drag-reduction, and enhanced heat and mass transfer [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The most common silicone is polydimethylsiloxane (PDMS). PDMS is biocompatible, water repellent, and flexible while being chemically/thermally stable [11, 12, 13, 14, 15, 16, 17, 18, 19]. The mechanical properties of PDMS can be tuned over many orders of magnitude and the synthesis is safe and mostly straightforward when using commercially available two-part systems, requiring no advanced training. Thus, PDMS is readily used in labs as a model system for soft materials.
Due to the chemical and mechanical properties, PDMS is rigorously used as a surface coating: The softness and flexibility of PDMS yield excellent lubrication properties [20, 21, 22] which can be utilized for joint lubrication [23] or drag reduction [24, 25]. The low (lateral) adhesion is exploited for applications in anti-icing [26, 8, 27, 28], anti-marine fouling [29], and anti-microbial formation [30, 31, 32, 33, 34]. Heat and mass transfer can be optimized as PDMS surfaces provide high nucleation rates while maintaining drop mobility [35, 36, 37]. PDMS is also used to interface biological systems, e.g., to control the growth and the size of cells [38, 39].
@@
To provide optimal surface functionality, understanding the wetting interactions in a given application is essential. Misinterpretation of these interactions can lead to misuse, malfunction, and deterioration of the surface. However, such understanding is challenging as the interactions are complex with strong coupling between the different components, and require consideration on many levels. During wetting, various forms of mass, momentum, and energy exchange take place on length scales ranging from the molecular structure of the polymer ($\mathrm{\AA}$) to the macroscopic drop ($\mathrm{cm}$) [41, 42, 43]. The associated timescales vary from mere picoseconds up to several days. The liquid drop can mix to some extent with the underlying polymer [44, 45, 46]. The sub-components that make up the polymer matrix of solid PDMS can undergo chemical reactions or phase changes. This alters the molecular structure of either or both the surface and the liquid [47, 41]. Liquid PDMS chains (e.g., oil) can engulf and cloak the drop [48, 49, 50]. The cloak changes the drop affinity to the surface and therefore affects the wettability. Interactions between the drop and surface are particularly high near the so-called ”three-phase-contact line” [51, 52]. At this location, the drop meets with the surface and the surrounding medium (a third fluid such as air). Here, the surface tension of the sessile drop exerts pulling stress on the surface. On many rigid surfaces such as metals or glass, the pulling stress is minor compared to the internal material stresses. PDMS, however, is soft, even when it comes as a rubber. Hence, the surface tension-induced pulling stress deforms the surfaces to an annular ”wetting ridge” [53, 54, 55, 56]. On very soft surfaces the wetting ridge grows large enough that it becomes visible by the bare eye.
@@
After establishing the structural differences in chain-chain and chain-surface organization on the different surfaces, we discuss their characteristic wetting behavior. We consider static and dynamic wetting - with a particular focus on the wetting ridge. This feature is ubiquitous amongst PDMS-based surfaces and the most pivotal element for wetting, drop mobility, and surface durability. Contact friction mostly dissipates in the ridge. Therefore it is critical for lubrication and drop sliding. However, it also accumulates surface material, and contacting objects may entrain surface material in the wetting ridge [63]. Eventually, this can cause surface deterioration and malfunction. We discuss secondary effects such as PDMS-drop cloaking which influences the wetting behavior and surface durability.
@@ II PDMS-Based Surface - Structure and Properties
The siloxane backbone is more flexible than a carbon backbone. This flexibility is attributed to the large $\mathrm{Si-O-Si}$ bond angle of $143^{\circ}-150^{\circ}$ [65, 66, 62] which separates the pendant dimethyl groups. In addition, the alternating oxygen atoms do not allow side groups, resulting in low rotational energy around the backbone ($3.3~{}\mathrm{kJ/mol}$) - e.g., lower than what is noted between carbon-carbon bonds with methylene $\mathrm{CH_{2}}$ configurations ($13.8~{}\mathrm{kJ/mol}$) that typically occur in comparable organic polymers, such as polyethylene glycol [67]. As a result, PDMS chains are highly flexible, and readily re-orientate to a minimal energy state, while possessing a low glass transition temperature ($\approx-150^{\circ}~{}\mathrm{C}$ [68]).
The organic content in PDMS is relatively low compared to other silicones due to the small organic methyl group ($\mathrm{CH_{3}}$, is the smallest organic group). The apolarity of the methyl group yields weak interactions with polar solvents such as water and alcohols [69, 70]. This is reflected in the low solubility which, in the case of water, is only $\approx 770~{}\mathrm{ppm}$ [71, 72, 73]. Vice versa, PDMS is even insoluble in water [29]. The poor interaction strength between PDMS and water, however, does not hamper the diffusive mobility of $\mathrm{H_{2}O}$ molecules in PDMS ($\approx 2\times 10^{-9}~{}\mathrm{m^{2}/s}$ [74, 73]) and is comparable with many other liquid-liquid diffusivities.
@@
PDMS chains can be adjusted in length (i.e., molecular weight), cross-linkage, and chemical/physical surface attachments (”grafting”). Such variations shape distinctive surface types and features including viscosity, elasticity, film thickness, and durability.
@@ II.1 Liquid-Infused Surfaces
The design of liquid-infused surfaces is to some extent nature-inspired: carnivorous plants such as the Nepenthes pitcher plant utilize a combination of leaf texture and lubrication (the plant uses water as a lubricant) to trap insects such as ants [85]. Liquid-infused surfaces comprise a micro- or nanostructured texture, infiltrated with liquid PDMS oil, also termed silicone oil. Typically, silicone oils are PDMS chains terminated with trimethyl groups, Fig 2a. The end groups are nonreactive; therefore PDMS remains liquid with a linear chain topology [13], Fig. 2a, left. The oil viscosity is tunable via the number of monomer units in the chain and ranges from $1~{}\mathrm{mPa~{}s}$ to $10~{}\mathrm{kPa~{}s}$ [86]. Short PDMS chains (”oligomers” [87]) with $n<100$ monomer units are typically very mobile with low viscosities of $\eta\leq 50~{}\mathrm{mPa~{}s}$. Due to the low surface tension of PDMS oils, almost all surfaces are ”siliconephilic”, enabling rapid spreading of oil into the porous, micro-/nanostructure textures [88, 89, 90]. Micro- and/or nanostructures are added to the surface to enable stronger oil retention by capillary forces [91]. These surfaces are summarized under the surface class named ”SLIPS” (slippery liquid-infused porous surface) or ”LIS” (lubricant-infused surface) [58, 57]. The oil provides lubrication, resulting in static friction forces orders of magnitude lower than dry friction. Most liquid drops slide off when tilting the surface by merely a few degrees. However, the infused oil can also be taken along, inducing degradation of the coating.
@@ II.2 Elastomeric Surfaces
The mixture of activated PDMS chains and cross-linker is coated on a surface by e.g., spin-coating, dip-coating, or blade-coating [101, 102, 103], before the reaction proceeds. Crosslinking reactions (independent of their formulation) will eventually arrive at the ”gel point”. The gel point is achieved when a percolated bulk network forms and the material loses fluidity [104]. Physically, this is an equilibrium state where the original PDMS liquid mixture becomes a soft solid (i.e., a gel). Three specific changes to physicochemical properties arise. i) Thermal rheology; as with other thermoset polymers, once a PDMS elastomer reaches the gel point, it can no longer reflow upon heating (unlike thermoplastics). This occurs above the glass transition temperature (or even melting point) of an equivalent linear (identical $\mathrm{M_{w}}$, uncrosslinked) PDMS chain. ii) Solvation; prior to gelation, the PDMS liquid mixture is readily soluble in certain solvents (e.g., hexane or toluene [69]). However, post-gelation, cross-linked PDMS will experience swelling instead. The cross-links preserve the integrity of the elastomeric polymer network. iii) Mechanical; passing the gel point induces an abrupt increase of the viscosity and the emergence of elasticity [105]. Depending on the cross-linker to oligomer ratio, the material obtains a stiffness between $1~{}\mathrm{kPa}$ and $10~{}\mathrm{MPa}$ [106, 107, 79, 108, 109].
@@ II.3 SOCAL Surfaces
The free chain ends of PDMS can be covalently anchored to a surface, creating a nanometer-thin coating [115, 116, 117, 118, 119], Fig. 1c. Most chains are anchored at one end while the other remains free. Few chains can be grafted on both ends, forming chain loops [120, 121]. The free-moving end is almost as flexible as liquid oligomers or oils and the surface interface is comparably slippery. Therefore, these surfaces are often called ”liquid-like” or ”SOCAL” (slippery, omniphobic, covalently attached, liquid) surfaces.
Surface anchoring is done either by a ”grafting-from” or a ”grafting-to” reaction. ”Grafting-from” implies that chain polymerization is initiated at the anchor site on the surface and proceeds by adding PDMS monomer units. In a ”grafting-to” reaction, an already polymerized chain binds to the surface [122, 115, 116]. ”Grafting-from” can be attained by e.g., highly reactive chloro-terminated siloxanes (monomers or oligomers). Long PDMS chains ($n>1000$) will experience ”grafting-to” anchoring. In both instances, they are reported to form covalent bonds and create molecularly thin layers [123].
Surface anchor sites are typically provided by hydroxyl groups, Fig. 2c. Anchor sites can be created using chemical treatment, e.g., oxygen plasma exposure or alkali activation. Many metals/metalloid oxides (e.g., $\mathrm{SiO_{2}}$, $\mathrm{TiO_{2}}$, $\mathrm{Al_{2}O_{3}}$, $\mathrm{NiO}$) naturally form hydroxyl groups under ambient water vapor exposure [124].
@@ III.1 Static Wetting Ridge
All surface types share the formation of an annular wetting ridge around the three-phase contact line [129, 130, 56, 131]. The wetting ridge is defined as the region that rises above the original unperturbed level of the surface. On liquid-infused surfaces, the wetting ridge consists of PDMS oil only, whereas on elastomers, it may consist of a mixture of crosslinked and liquid PDMS, depending on the swelling ratio. On SOCAL surfaces, the wetting ridge will encompass the local stretching of the grafted polymer chains. As PDMS chains are organized differently on the liquid-infused, elastomer, and SOCAL surfaces, the material forces in the wetting ridge differ, as well. Thus, each surface type forms a characteristic wetting ridge with a specific geometry. The wetting ridge significantly contributes to the overall drop morphology and the wetting dynamics.
Figure 3: Different manifestations of the wetting ridge on different PDMS surfaces. a) Liquid wetting ridge on a liquid-infused surface. Neumann balance at the three-phase contact line (box) and the apparent contact angle $\theta_{app}$. The right half illustrates the starved limit and the pseudo-Young’s balance. Macroscopically, the vector center of the surface tensions aligns with the baseline of the substrate. The image on the right shows a liquid wetting ridge, taken with a fluorescence confocal microscope, adapted from [92]. b) Elastic wetting ridge on an elastomeric surface. Bulk elasticity creates a compliance force $f_{el}\propto\mu\delta_{e,\mathrm{I}}$, resulting in wetting ridges smaller than the liquid ones. The right half shows material dimpling around the wetting ridge when the surface coating is very thin. The right image shows an elastic wetting ridge, taken with an x-ray microscope on elastomeric PDMS ($3~{}\mathrm{kPa}$), adapted from [132]. c) Molecular wetting ridge on a SOCAL surface. The entropic force of the PDMS chain $f_{m}$ limits the surface tension-induced chain stretch. Right half shows surface swelling when liquid-PDMS affinity/solubility is high. The right image shows a molecular wetting ridge, computed with coarse-grained molecular dynamics simulation, adapted from [133].
@@ III.1.1 Liquid
In addition to the hydrophobic nature of PDMS oil, the impregnated oil acts as a lubricant. The film lubrication reduces the (static) friction of contacting objects (both solids and liquids) on the surface immensely. Low friction typically leads to excellent self-cleaning properties [57]. Furthermore, lubricant allows surfaces to ”self-heal” from abrasion damages by the replenishment and re-flow of oil. However, once the PDMS oil is partially depleted from the surface, contacting drops are no longer lubricated and lose their mobility [134]. Maintaining the PDMS oil on the surface is, hence, the most important priority for the design of liquid-infused surfaces [91, 135, 28].
@@
For drops smaller than the capillary length, gravity can be neglected and capillary interactions between the lubricant and the water drop solely govern the morphology of the wetting ridge. In contrast to wetting on inert rigid surfaces, a single angle, such as the Young angle [136], is insufficient for this wetting configuration [137]. On soft surfaces, the surface material (the lubricant) adapts to the wetting drop as the interfacial tension of the drop pulls on the surface. The surface becomes curved in the vicinity of the wetting ridge. In equilibrium, the (liquid) wetting ridge is in a so-called ”Neumann configuration”, which implies a balance of the three fluid interfacial tensions [$\vec{\gamma}_{w}+\vec{\gamma}_{o}+\vec{\gamma}_{ow}=0,$ subscripts correspond oil-ambient, water-ambient, and oil-water], Fig. 3a left box. This tension balance yields a closed triangle, characterized by the associated three Neumann angles $\vartheta_{1}$, $\vartheta_{2}$, and $\vartheta_{3}$ [138]. The angles are related to the interfacial tensions as
@@
For water and PDMS oil, $\gamma_{w}>\gamma_{o}+\gamma_{ow}$, hence no Neumann configuration exists, and we expect the oil to fully spread around the drop (cf. Sec. Free Chain Cloaking Layer). In other situations, the height of the wetting ridge (or the location of the three-phase contact line) keeps growing until the lubricant pressure in the wetting ridge matches the lubricant pressure in the solid textures. In equilibrium, each of the fluid interfaces assumes a minimal energy surface [139] where the mean curvature is constant [140, 141, 142]. The oil interface deforms significantly in the surroundings of the wetting ridge but straightens, afar. The ”catenoid” surface approximates all these properties. This is possible because the surface curvature in the two principal normal directions has equal magnitude but opposite signs, such that they perfectly cancel out.
While the classical Young equation/angle fails for lubricant-infused surfaces, a similar angle can be recovered in the limit when the wetting ridge is small compared to the size of the drop, Fig. 3a center. This limit typically prevails in the ”starved limit” when the infused PDMS oil is scarce. The resulting ”pseudo Young angle” $\theta_{app}^{S}$ [143] is
@@
The effective surface tensions of the solid have contributions from both, texture ($\Upsilon_{s}$ and $\Upsilon_{sl}$) and PDMS oil ($\gamma_{o}$ and $\gamma_{ow}$), i.e., $\Upsilon_{s}^{\mathrm{eff}}=\Upsilon_{s}\varphi+\gamma_{o}(1-\varphi)$ and $\Upsilon_{sl}^{\mathrm{eff}}=\Upsilon_{sl}\varphi+\gamma_{ow}(1-\varphi)$, where $\varphi$ is the texture fraction of the solid surface.
Moreover, in the starved limit, the shape of the wetting ridge is not only defined by capillarity. Disjoining pressures must also be considered when the oil film becomes very thin ($\lessapprox 100~{}\mathrm{nm}$) because the oil-ambient interface starts to interact with the oil-solid interface. Such interactions are characteristic of very thin films and are summarised in the ”disjoining pressure” framework [144]. The interactions take place on a molecular level and can stem from a combination of van der Waals interactions, electrostatic attraction or repulsion, and steric forces [145]. For static PDMS oil films, the interactions are usually purely dispersive, and the disjoining pressure is
@@
When the surface holds sufficient PDMS oil, the contact angle defined in Eq. (2) breaks down because the assumption of a small wetting ridge no longer holds. The interface of the ridge now contributes to the total free energy. It becomes challenging to define a baseline from which to measure the (pseudo) angle, as the wetting ridge conceals the footprint of the drop. In practice, apparent contact angles $\theta_{app}$ are utilized, which are defined as the angle of the drop at the Neumann point to the horizontal plane, Fig. 3a left. The size of the large ridge $\delta_{l}$ and the size of the drop $r$ are related to the difference between $\theta_{app}$ and $\theta_{app}^{S}$ [149] per
$\delta_{l}=r\left(\cos{\theta_{app}}-\cos{\theta_{app}^{S}}\right).$ (4)

### v2

@@ I Introduction
Silicones are abundant materials with global production margins exceeding a million tonnes annually. As a surface coating, the material is widely used in numerous industries, ranging from the medical and energy sectors to the personal care and automotive/aerospace industries. The low surface energy of silicones promotes low friction and adhesion with contacting solids and liquids, providing desirable properties for self-cleaning, lubrication, anti-icing/-biofouling, drag-reduction, and enhanced heat and mass transfer [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The most common silicone is polydimethylsiloxane (PDMS). PDMS is biocompatible, water repellent, and flexible while being chemically/thermally stable [11, 12, 13, 14, 15, 16, 17, 18, 19]. The chemical inertness, low miscibility in water, and non-toxicity make PDMS mostly environmentally harmless [20]. Recent bacterial studies showed a biodegradability of the polymer [21], enabling a sustainable product life cycle. The mechanical properties of PDMS can be tuned over many orders of magnitude and the synthesis is safe and mostly straightforward when using commercially available two-part systems, requiring no advanced training. Thus, PDMS is readily used in labs as a model system for soft materials.
Due to the chemical and mechanical properties, PDMS is used as a surface coating: The softness and flexibility of PDMS yield excellent lubrication properties [22, 23, 24] which can be utilized for joint lubrication [25] or drag reduction [26, 27]. The low (lateral) adhesion is exploited for applications in anti-icing [28, 8, 29, 30], anti-marine fouling [31], and anti-microbial formation [32, 33, 34, 35, 36]. Heat and mass transfer can be optimized as PDMS surfaces provide high nucleation rates while maintaining drop mobility [37, 38, 39]. PDMS is also used to interface biological systems, e.g., to control the growth and the size of cells [40, 41].
@@
To provide optimal surface functionality, understanding the wetting interactions in a given application is essential. Misinterpretation of these interactions can lead to misuse, malfunction, and deterioration of the surface. However, such understanding is challenging as the interactions are complex with strong coupling between the different components, and require consideration on many levels. During wetting, various forms of mass, momentum, and energy exchange take place on length scales ranging from the molecular structure of the polymer (Å) to the macroscopic drop ($\mathrm{cm}$) [43, 44, 45]. The associated timescales vary from mere picoseconds up to several days. The liquid drop can mix to some extent with the underlying polymer [46, 47, 48]. The sub-components that make up the polymer matrix of solid PDMS can undergo chemical reactions or phase changes. This alters the molecular structure of either or both the surface and the liquid [49, 43]. Liquid PDMS chains (e.g., oil) can engulf and cloak the drop [50, 51, 52]. The cloak changes the drop affinity to the surface and therefore affects the wettability. Interactions between the drop and surface are particularly high near the so-called “three-phase-contact line” [53, 54]. At this location, the drop meets with the surface and the surrounding medium (a third fluid such as air). Here, the surface tension of the sessile drop exerts pulling stress on the surface. On many rigid surfaces such as metals or glass, the pulling stress is minor compared to the internal material stresses. PDMS, however, is soft, even when it comes as a rubber. Hence, the surface tension-induced pulling stress deforms the surfaces to an annular “wetting ridge” [55, 56, 57, 58]. On very soft surfaces the wetting ridge grows large enough that it becomes visible by the bare eye.
@@
After establishing the structural differences in chain-chain and chain-surface organization on the different surfaces, we discuss their characteristic wetting behavior. We consider static and dynamic wetting - with a particular focus on the wetting ridge. This feature is ubiquitous amongst PDMS-based surfaces and the most pivotal element for wetting [65, 66, 67, 68, 69, 70, 71], drop mobility [70, 72, 73, 65, 74, 75, 76], and surface durability [77, 78, 79, 80, 81, 46, 82]. Contact friction mostly dissipates in the ridge [83, 74, 73, 84, 85]. Therefore it is critical for lubrication and drop sliding. However, it also accumulates surface material, and contacting objects may entrain surface material in the wetting ridge [86]. Eventually, this can cause surface deterioration and malfunction. We discuss secondary effects such as PDMS-drop cloaking [87, 88, 89] which influences the wetting behavior and surface durability.
@@ II PDMS-Based Surface - Structure and Properties
The siloxane backbone is more flexible than a carbon backbone. This flexibility is attributed to the large $\mathrm{Si-O-Si}$ bond angle of $143^{\circ}-150^{\circ}$ [91, 92, 64] which separates the pendant dimethyl groups. In addition, the alternating oxygen atoms do not allow side groups, resulting in low rotational energy around the backbone ($3.3~{}\mathrm{kJ/mol}$) - e.g., lower than what is noted between carbon-carbon bonds with methylene $\mathrm{CH_{2}}$ configurations ($13.8~{}\mathrm{kJ/mol}$) that typically occur in comparable organic polymers, such as polyethylene glycol [93]. As a result, PDMS chains are highly flexible, and readily re-orientate to a minimal energy state, while possessing a low glass transition temperature ($\approx-150^{\circ}~{}\mathrm{C}$)[94].
The organic content in PDMS is relatively low compared to other silicones due to the small organic methyl group ($\mathrm{CH_{3}}$, is the smallest organic group). The apolarity of the methyl group yields weak interactions with polar solvents such as water and alcohols [95, 96]. This is reflected in the low solubility which, in the case of water, is only $\approx 770~{}\mathrm{ppm}$ [97, 98, 99]. Vice versa, PDMS is even insoluble in water [31]. The poor interaction strength between PDMS and water, however, does not hamper the diffusive mobility of $\mathrm{H_{2}O}$ molecules in PDMS ($\approx 2\times 10^{-9}~{}\mathrm{m^{2}/s}$)[100, 99] and is comparable with many other liquid-liquid diffusivities.
@@
PDMS chains can be adjusted in length (i.e., molecular weight), cross-linkage, and chemical/physical surface attachments (“grafting”). Such variations shape distinctive surface types and features including viscosity, elasticity, film thickness, and durability.
@@ II.1 Liquid-Infused Surfaces
The design of liquid-infused surfaces is to some extent nature-inspired: carnivorous plants such as the Nepenthes pitcher plant utilize a combination of leaf texture and lubrication (the plant uses water as a lubricant) to trap insects such as ants [111]. Liquid-infused surfaces comprise a micro- or nanostructured texture, infiltrated with liquid PDMS oil, also termed silicone oil. Typically, silicone oils are PDMS chains terminated with trimethyl groups, Fig 2a. The end groups are nonreactive; therefore PDMS remains liquid with a linear chain topology [13], Fig. 2a, left. The oil viscosity is tunable via the number of monomer units in the chain and ranges from $1~{}\mathrm{mPa~{}s}$ to $10~{}\mathrm{kPa~{}s}$ [112]. Short PDMS chains (“oligomers” [113]) with $n<100$ monomer units are typically very mobile with low viscosities of $\eta\leq 50~{}\mathrm{mPa~{}s}$. Due to the low surface tension of PDMS oils, almost all surfaces are “siliconephilic”, enabling rapid spreading of oil into the porous, micro-/nanostructure textures [114, 115, 116]. Micro- and/or nanostructures are added to the surface to enable stronger oil retention by capillary forces [117]. These surfaces are summarized under the surface class named “SLIPS” (slippery liquid-infused porous surface) or “LIS” (lubricant-infused surface) [60, 59]. The oil provides lubrication, resulting in static friction forces orders of magnitude lower than dry friction. Most liquid drops slide off when tilting the surface by merely a few degrees. However, the infused oil can also be taken along, inducing degradation of the coating.
@@ II.2 Elastomeric Surfaces
The mixture of activated PDMS chains and cross-linker is coated on a surface by e.g., spin-coating, dip-coating, or blade-coating [127, 128, 129], before the reaction proceeds. Crosslinking reactions (independent of their formulation) will eventually arrive at the “gel point”. The gel point is achieved when a percolated bulk network forms and the material loses fluidity [130]. Physically, this is an equilibrium state where the original PDMS liquid mixture becomes a soft solid (i.e., a gel). Three specific changes to physicochemical properties arise. i) Thermal rheology; as with other thermoset polymers, once a PDMS elastomer reaches the gel point, it can no longer reflow upon heating (unlike thermoplastics). This occurs above the glass transition temperature (or even melting point) of an equivalent linear (identical $\mathrm{M_{w}}$, uncrosslinked) PDMS chain. ii) Solvation; prior to gelation, the PDMS liquid mixture is readily soluble in certain solvents (e.g., hexane or toluene [95]). However, post-gelation, cross-linked PDMS will experience swelling instead. The cross-links preserve the integrity of the elastomeric polymer network. iii) Mechanical; passing the gel point induces an abrupt increase of the viscosity and the emergence of elasticity [131]. Depending on the cross-linker to oligomer ratio, the material obtains a stiffness between $1~{}\mathrm{kPa}$ and $10~{}\mathrm{MPa}$ [132, 133, 105, 134, 135].
@@ II.3 SOCAL Surfaces
The free chain ends of PDMS can be covalently anchored to a surface, creating a nanometer-thin coating [137, 138, 139, 140, 141], Fig. 1c. Most chains are anchored at one end while the other remains free. Few chains can be grafted on both ends, forming chain loops [142, 143]. The free-moving end is almost as flexible as liquid oligomers or oils and the surface interface is comparably slippery. Therefore, these surfaces are often called “liquid-like” or “SOCAL” (slippery, omniphobic, covalently attached, liquid) surfaces.
Surface anchoring is done either by a “grafting-from” or a “grafting-to” reaction. “Grafting-from” implies that chain polymerization is initiated at the anchor site on the surface and proceeds by adding PDMS monomer units. In a “grafting-to” reaction, an already polymerized chain binds to the surface [144, 145, 137, 138]. “Grafting-from” can be attained by e.g., highly reactive chloro-terminated siloxanes (monomers or oligomers). Long PDMS chains ($n>1000$) will experience “grafting-to” anchoring. In both instances, they are reported to form covalent bonds and create molecularly thin layers [146].
Surface anchor sites are typically provided by hydroxyl groups, Fig. 2c. Anchor sites can be created using chemical treatment, e.g., oxygen plasma exposure or alkali activation [147]. Many metals/metalloid oxides (e.g., $\mathrm{SiO_{2}}$, $\mathrm{TiO_{2}}$, $\mathrm{Al_{2}O_{3}}$, $\mathrm{NiO}$) naturally form hydroxyl groups under ambient water vapor exposure [148]. Water dissociates on surfaces if acids (e.g. hydrofluoric acid) or alkali (e.g. sodium hydroxide) are used. Spontaneous dissociation of water on metal surfaces has also been reported [149, 150], alongside the spontaneous scission of covalent bonds in long PDMS chains [137].
@@ III.1 Static Wetting Ridge
All surface types share the formation of an annular wetting ridge around the three-phase contact line [155, 156, 58, 157]. The wetting ridge is defined as the region that rises above the original unperturbed level of the surface. On liquid-infused surfaces, the wetting ridge consists of PDMS oil only [158, 118, 159], whereas on elastomers, it may consist of a mixture of crosslinked and liquid PDMS, depending on the swelling ratio [79, 81]. On SOCAL surfaces, the wetting ridge will encompass the local stretching of the grafted polymer chains [160, 161]. As PDMS chains are organized differently on the liquid-infused, elastomer, and SOCAL surfaces, the material forces in the wetting ridge differ, as well. Thus, each surface type forms a characteristic wetting ridge with a specific geometry. The wetting ridge significantly contributes to the overall drop morphology and the wetting dynamics.
@@ III.1.1 Liquid
In addition to the hydrophobic nature of PDMS oil, the impregnated oil acts as a lubricant. The film lubrication reduces the (static) friction of contacting objects (both solids and liquids) on the surface immensely. Low friction typically leads to excellent self-cleaning properties [59]. Furthermore, lubricant allows surfaces to “self-heal” from abrasion damages by the replenishment and re-flow of oil. However, once the PDMS oil is partially depleted from the surface, contacting drops are no longer lubricated and lose their mobility [162]. Maintaining the PDMS oil on the surface is, hence, the most important priority for the design of liquid-infused surfaces [117, 163, 30].
@@
For drops smaller than the capillary length, gravity can be neglected and capillary interactions between the lubricant and the water drop solely govern the morphology of the wetting ridge. In contrast to wetting on inert rigid surfaces, a single angle, such as the Young angle [164], is insufficient for this wetting configuration [165]. On soft surfaces, the surface material (the lubricant) adapts to the wetting drop as the interfacial tension of the drop pulls on the surface. The surface becomes curved in the vicinity of the wetting ridge. In equilibrium, the (liquid) wetting ridge is in a so-called “Neumann configuration”, which implies a balance of the three fluid interfacial tensions [$\vec{\gamma}_{w}+\vec{\gamma}_{o}+\vec{\gamma}_{ow}=0,$ subscripts correspond water-ambient, oil-ambient, and oil-water], Fig. 3a left box. This tension balance yields a closed triangle, characterized by the threeassociated Neumann angles $\vartheta_{1}$, $\vartheta_{2}$, and $\vartheta_{3}$ [166]. The angles are related to the interfacial tensions as
@@
Drops of organic liquids (e.g., perflourinated oils) fulfill this condition [80]. Due to the high surface tension of water, however, $\gamma_{w}>\gamma_{o}+\gamma_{ow}$, and hence no Neumann configuration exists., and we expect the oil to fully spread around the drop (cf. Sec. Free Chain Cloaking Layer). In other situations, the height of the wetting ridge (or the location of the three-phase contact line) keeps growing until the lubricant pressure in the wetting ridge matches the lubricant pressure in the solid textures. In equilibrium, each of the fluid interfaces assumes a minimal energy surface [167] where the mean curvature is constant [168, 169, 170]. The oil interface deforms significantly in the surroundings of the wetting ridge but straightens, afar. The “catenoid” surface approximates all these properties. This is possible because the surface curvature in the two principal normal directions has equal magnitude but opposite signs, such that they perfectly cancel out.
While the classical Young equation/angle fails for lubricant-infused surfaces, a similar angle can be recovered in the limit when the wetting ridge is small compared to the size of the drop, Fig. 3a center. This limit typically prevails in the “starved limit” when the infused PDMS oil is scarce. The resulting “pseudo Young angle” $\theta_{app}^{S}$ [171] is
@@
The surface tension of the solid texture and the PDMS oil are $\Upsilon_{s}$ and $\Upsilon_{sl}$, respectively. The solid texture takes up a fraction $\varphi$ of the overall surface. Thus, effective surface tension can be introduced, i.e., $\Upsilon_{s}^{\mathrm{eff}}=\Upsilon_{s}\varphi+\gamma_{o}(1-\varphi)$ for the solid texture and $\Upsilon_{sl}^{\mathrm{eff}}=\Upsilon_{sl}\varphi+\gamma_{ow}(1-\varphi)$ for the surface-bound PDMS.
Moreover, in the starved limit, the shape of the wetting ridge is not only defined by capillarity. Disjoining pressures must also be considered when the oil film becomes very thin ($<\approx 100~{}\mathrm{nm}$) because the oil-ambient interface starts to interact with the oil-solid interface [172]. Such interactions are characteristic of very thin films and are summarised in the “disjoining pressure” framework [172]. The interactions take place on a molecular level and can stem from a combination of van der Waals interactions, electrostatic attraction or repulsion, and steric forces [173]. For static PDMS oil films, the interactions are usually purely dispersive, and the disjoining pressure is
@@
When the surface holds sufficient PDMS oil, the contact angle defined in Eq. (2) breaks down because the assumption of a small wetting ridge no longer holds. The interface of the ridge now contributes to the total free energy. It becomes challenging to define a baseline from which to measure the (pseudo) angle, as the wetting ridge conceals the footprint of the drop. In practice, apparent contact angles $\theta_{app}$ are utilized, which are defined as the angle of the drop at the Neumann point to the horizontal plane, Fig. 3a left. The size of the large ridge $\lambda_{l}$ and the size of the drop $r$ are related to the difference between $\theta_{app}$ and $\theta_{app}^{S}$ [67] per
$\lambda_{l}=r\left(\cos{\theta_{app}}-\cos{\theta_{app}^{S}}\right).$ (4)